An electric train start from a station and maintained an acceleration of 0.9m/s2 for 20 seconds, if it then travel for 80 seconds with uniform velocity and finally has a uniform retardation which brings it to rest in a further 10 seconds. calculate the retardation and the total distance covered.


1
Expert's answer
2021-03-08T03:09:04-0500

Speed after 20 seconds:


"v=at"

Distance covered in 20 s:


"d=vt\/2"

Distance covered in 80 seconds of uniform motion:


"x=vT=atT"

Retardation:


"a_r=v/\tau=at/\tau=1.8\text{ m/s}^2."

Distance in retardation:


"r=v\tau/2=at\tau/2"

Total distance:


"X=d+x+r=\\
=at(0.5+T+0.5\tau)=\\
=0.9·20(0.5+80+0.5·10)=1539\text{ m}."